15. Lara has been invited to a picnic on Saturday. The picnic will occur, rain of shine. If it rains, there is a 20% probability that Lara will decide to go, but if it's sunny, there is an 80% chance of rain. What is the probability that Lara will attend the picnic?
16. Three cards are dealt at random from a standard deck of 52 cards. What is the probability that:
(a) The first card is a Jack, the second card is a Queen, and the third card is a King?
(b) All three cards are hearts?
(c) The first card is a 4, the second card is a club, and the third card is a 2?

Answer by edjones(8007)

16.
a) 4/52 * 4/51 * 4/50 = .00048
.
b) 13/52 * 12/52 * 11/52 = .0122
.
c) 4/52 * ((3(13/51) + 12/51)/4)* ((12(4/50) + 3/50) / 13) = .0015 The 1st card could have been the club 4 and the last, the club 2 might have been selected as a club.
.
Ed